OpenStudy (anonymous):

Solve the system, using substitution. Show work please! thanks! (a,b). a + b = 4 and 2a + 3b = 11

OpenStudy (anonymous):

In your first equation, solve for either variable. Then substitute what you get from that, into the second equation. Should become obvious from there.

OpenStudy (anonymous):

a=4-b substitute a in 2nd equation

OpenStudy (anonymous):

Yes what they said... make your other equation 2(4-b)+3b=11

OpenStudy (anonymous):

So that will become 8-2b+3b=11

OpenStudy (anonymous):

Simplify that into 8+b=11

OpenStudy (anonymous):

that turns into b=3

OpenStudy (anonymous):

Now take your value for b and substitute it into the equation

OpenStudy (anonymous):

2(4-b)+3b=11 8-2b+3b=11 8+b=11 b=3

OpenStudy (anonymous):

a+(3)=4

OpenStudy (anonymous):

a=1 so b=3 and a=1